Q4. ABC Ltd. manufactures luxury golf bags in their two plants located in Augusta and Tupper Lake Warchouse facilities are located in Albany and Portsmouth. Distributors are located in Boston, New York and Philadelphia. The Augusta plant has a capacity of 300 units per month, and the Tupper Lake plant has a capacity of 100 units per month. Boston has a demand of 150 units per month, New York has a demand of 100 units per month, and Philadelphia has a demand of 150 units per month. The unit transportation costs (in dollars) for shipments from the two plants to the two warehouses are presented in Table 3 and those from the two warchouses to the three distributors are presented in Table 4 Table 3. Plant to Warehouse Information Warehouse Albany Portsmouth Piant Augusta Tupper Lake 3 4 Table 4. Warehouse to Distributor Information Distributor Boston New York Philadelphia Warehouse Albany Portsmouth 10 12 Marks a) Draw the network representation of the above Transhipment problem b) Formulate a linear programming model to determine the optimal quantity of bags to be transported from plants to warehouses, and from warehouses to distributors so that total transportation cost is minimized as well as total demand at distributors is satisfied. 5 Marks) e) An internationa! golftourzament is scheduled in New York which will increase the demanl of luxury golf bags in New York by 50 units next month. Formulate a feasible linear programming model by incorporating this increment in demand. 3 Marks) Q5. Quality Air Conditioning manufactures three types of home air conditioners: an economy model, a el The nrefits per unit are $63, $95, and $135, respectively. The
Answer #1

a) Network represetnation is as follows:

Plant Warehouse Distributor Boston-150 New York |-100 Philadelphia -150 300 Augusta Albany 100 Tupper Lake Portsmouth

b) LP model

Decision variable:

Xij = Quantity to be shipped from plant i to Warehousej

Xjk = Quantity to be shipped from Warehouse j to Distributor k

where  i\in \left \{ 1,2 \right \}, j\in\left \{ 3,4 \right \}, k\in\left \{ 5,6,7 \right \}

Objective:

Min: \sum_{i=1}^{2}\sum_{j=3}^{4}X_{ij}*C_{ij} + \sum_{j=3}^{4}\sum_{k=5}^{7}X_{jk}*C_{jk}

s.t.

X13 + X14 = 300

X23 + X24 = 100

X13 + X23 = X35 + X36 + X37

X14 + X24 = X45 + X46 + X47

X35 + X45 = 150

X36 + X46 = 100

X37 + X47 = 150

Xij >= 0

Solution using Excel Solver follows:

B24 Q fx -SUMPRODUCT(B3:C4,B14:C15)+SUMPRODUCT(B8:D9,B20:D21) 1 Table 3: Plant to Warehouse 2 Plant Warehouse 3 Augusta 4 Tup

Formula:

B16 =SUM(B14:B15) copy to B16:C16, B22:D22

D14 =SUM(B14:C14) copy to D14:D15

E20 =SUM(B20:D20) copy to E20:E21

B24 =SUMPRODUCT(B3:C4,B14:C15)+SUMPRODUCT(B8:D9,B20:D21)

c) The revised model is as follows:

Objective:

Min: \sum_{i=1}^{2}\sum_{j=3}^{4}X_{ij}*C_{ij} + \sum_{j=3}^{4}\sum_{k=5}^{7}X_{jk}*C_{jk}

s.t.

X13 + X14 = 300

X23 + X24 = 100

X13 + X23 = X35 + X36 + X37

X14 + X24 = X45 + X46 + X47

X35 + X45 <= 150

X36 + X46 <= 150

X37 + X47 <= 150

Xij >= 0

a fx SUMPRODUCT(B3:C4,B14:C15)+SUMPRODUCT(B8:D9,B20:D21) B24 1 Table 3: Plant to Warehouse 2 Plant Warehouse 3 Augusta 4 Tupp